OpenStudy (anonymous):

Find y' assuming that the equation determines a differentiable function f such that y= f(x) sin^2 3y=x+y-1

OpenStudy (amistre64):

\(sin^2(3y)\) ??

OpenStudy (anonymous):

its implicit differentiation

OpenStudy (amistre64):

yeah; but whats the equation?

OpenStudy (anonymous):

yes amistre that is correct

OpenStudy (anonymous):

its kind of assumed its sin^2 (3y) lol

OpenStudy (anonymous):

its unlikely to be sin^2(3) y

OpenStudy (amistre64):

\[3y' 2\cos(3y)=1+y'\] \[6y'\cos(3y)-y'=1\] \[y'(6\cos(3y)-1)=1\] \[y' = \frac{1}{6\cos(3y)}\]

OpenStudy (amistre64):

thats wrong ....

OpenStudy (anonymous):

?

OpenStudy (amistre64):

i did sin^2(3y) as tho it was simply sin(3y)

OpenStudy (anonymous):

ohh yeh, I see

OpenStudy (anonymous):

alot of chain rules in that question

OpenStudy (anonymous):

its only once elecengineer

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

\[3y'.2(\sin(3y)).cos(3y)=1+y'\] \[y'.6\sin(3y)\cos(3y)-y'= 1\] \[y'(6\sin(3y)\cos(3y)-1)= 1\] \[y' = \frac{1}{6\sin(3y)\cos(3y)-1}\] maybe

OpenStudy (amistre64):

3 times actually

OpenStudy (amistre64):

\[u^2.sin(v).3y\]

OpenStudy (anonymous):

also, you couls apply double angle formula for sin to simplify the bottom :P

OpenStudy (amistre64):

yup

OpenStudy (anonymous):

im trying to figure out what you did amistre

OpenStudy (anonymous):

consider the LHS first, the derivative of the RHS is easy as

OpenStudy (anonymous):

so ( sin(3y) )^2 apply the chain rule , bring the power down in front , leave the inside of the bracket alone , reduce the power by one , then multiplt by the derivative of the inside

OpenStudy (anonymous):

so d/dx [ (sin(3y) ) ^2 ] = 2 sin(3y) [d/dx ( sin(3y) ) ]

OpenStudy (anonymous):

= 2 sin(3y) 3y' cos(3y)

OpenStudy (amistre64):

say you wanna derive: u^2; thats simple right? but: u = sin(v) and v = 3y

OpenStudy (anonymous):

when you differentiate sin(3y) you take the derivative of the inside with respect to x, and then change the sin function to a cos function

OpenStudy (anonymous):

the derivative of 3y with respect to x is 3 (dy/dx) = 3y'

OpenStudy (anonymous):

i get the 2(sin3y)(cos y) but not the 3y'

OpenStudy (amistre64):

\[{du\over dx}={du\over dv}{dv\over dy}{dy\over dx}\] \[\frac{d(sin^2(3y))}{dx}=2sin(3y)*cos(3y)*3y'\]

OpenStudy (amistre64):

you dont understand why we keep something hta tyou were taught to throw out

OpenStudy (anonymous):

:S?

OpenStudy (amistre64):

\[y = 2x^3 \rightarrow y' = 6x^2 x'\]

OpenStudy (anonymous):

oh...

OpenStudy (amistre64):

what did the \(y\) derive to? \(y'\) right?

OpenStudy (amistre64):

3y derives to: \(3 y'\) \(5y^2\) derives to \(10y .y'\)

OpenStudy (anonymous):

so the derivative of sin 3y is 3 sin 3y then 3y' cos 3y?

OpenStudy (amistre64):

you are used to throwing out the x' bit; but that is only because \(dx\over dx\)=1

OpenStudy (anonymous):

i know that ist the trig function thats throwing me off

OpenStudy (amistre64):

sin(3y) derives to 3y' cos(3y)

OpenStudy (anonymous):

yes...thats what confused me or should i say confuses

OpenStudy (amistre64):

the sin to cos? or the innards to the outside?

OpenStudy (anonymous):

the innards the 3y inside the sin 3y

OpenStudy (amistre64):

the innards have a controling part in the function and have to be accounted for. They are the driving force that produces the output for the sin function

OpenStudy (amistre64):

the 'chain' rule is just accounting for all the inputs and outputs thru the equation that eah have a controling part

OpenStudy (anonymous):

ok i guess i have to practice some more examples of them

OpenStudy (amistre64):

try it on simple ones first: like, compound functions

OpenStudy (amistre64):

interactmath.com can help; its a free practice math site

OpenStudy (anonymous):

THANKS AGAIN SO MUCH AMISTRE AND ELECENGINEER

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

can you help me with the question i posted earlier today